# Relativistic quantum Otto engine: Instant work extraction from a quantum field

### 07 Dec 2023 v1

@@ 1 Introduction
While quantum heat engines have been extensively studied within the domain of quantum mechanics, their exploration in the context of quantum field theory remains relatively unexplored. Recent investigations UnruhOttoEngine ; Finn.UnruhOtto have delved into QOEs where the working substance is an Unruh-DeWitt (UDW) particle detector Unruh1979evaporation ; DeWitt1979 interacting with quantum scalar and spinor fields. In their QOE, the field is assumed to be in the Minkowski vacuum state, and work is extracted from a thermal bath induced by the Unruh effect Unruh1979evaporation experienced by a linearly accelerating detector. Such a QOE driven by the Unruh effect is particularly called the Unruh QOE. See also Xu.UnruhOtto.degenerate ; Kane.entangled.Unruh.Otto ; Barman.entangled.UnruhOtto ; Mukherjee.UnruhOtto.boundary for subsequent studies. The Unruh QOE concept extends to scenarios where a UDW detector follows an arbitrary timelike trajectory in curved spacetime Gallock2023Otto , which we will refer to as the relativistic QOE.
@@ 3.1 Algebraic quantum field theory
The CCR algebra $\mathcal{A}(\mathcal{M})$ is generated by smeared field $\phi(f)$ obeying the following relations khavkine2015algebraic , which we refer to the Klein-Gordon relations:
@@
Let us define a ***-representation of $\mathcal{A}(\mathcal{M})$ as a triple $(\mathcal{H},\mathcal{D},\pi)$, where $\mathcal{H}$ is Hilbert space, $\mathcal{D}\subset\mathcal{H}$ is a dense subspace of $\mathcal{H}$, and $\pi$ (also refer to as a representation) is a linear homomorphism,
@@ 5.2 Observations and remarks
    The positive work condition (55) implies that if the detector is unable to self-communicate, then $E(\mathfrak{f}_{1},\mathfrak{f}_{2})=0$ leading to $\braket{W_{\text{ext}}}=0$. Thus, self-communication is necessary to extract positive work for a delta-coupled UDW detector. It should be noted that merely receiving a signal from the past does not guarantee work extraction since the sign in (55) might be positive even when $E(\mathfrak{f}_{1},\mathfrak{f}_{2})\neq 0$. We will address the mechanism for the work extraction at the end of this subsection.
@@
The ranges of $p_{2}$ in (58) is due to the fact that $\nu_{\mathfrak{f}_{1}}\nu_{\mathfrak{f}_{2}}\alpha\in(0,1]$, as given by (51).
@@ 5.3 Example: An inertial detector extracts work from the Minkowski vacuum
Figure 5: The extracted work $\braket{W_{\text{ext}}}\sigma$ in units of the characteristic width $\sigma$ of the Gaussian smearing function. (a) Work as a function of the second interaction time $\tau_{2}/\sigma$ when $\tau_{1}/\sigma=0$, $\Omega_{1}\sigma=1$, $\Omega_{2}\sigma=3$, $\tilde{\lambda}_{1}/\sigma=100$, and $\tilde{\lambda}_{2}/\sigma=1$. Negative $\braket{W_{\text{ext}}}\sigma$ corresponds to work done on the detetor. (b) Work as a function of the coupling constants. Here, we choose $\tau_{1}/\sigma=0$, $\Omega_{1}\sigma=1$, $\Omega_{2}\sigma=3$, and $\tau_{2}/\sigma=1.5$.
@@
After substituting these expressions into the extracted work (54), the results are shown in Figure 5. In Figure 5(a), we depict the dependence of the second interaction’s timing, $\tau_{2}/\sigma$, in units of the characteristic Gaussian width $\sigma$, given $\tau_{1}/\sigma=0$, $\Omega_{1}\sigma=1$, $\Omega_{2}\sigma=3$, $\tilde{\lambda}_{1}/\sigma=100$, and $\tilde{\lambda}_{2}/\sigma=1$. As discussed in Section 5.2, the detector can extract positive work from the vacuum state of the field instantaneously as long as the first interaction affects the second one. In fact, we have $\braket{W_{\text{ext}}}\sigma=0$ once the detector cannot receive the signal from the first interaction (i.e., when $\tau_{2}/\sigma>4$).
@@ Appendix A Second interaction
$\nu_{\mathfrak{f}}$ $=\omega_{\mu}(W_{2\mathfrak{f}})=e^{-2\mu(E\mathfrak{f},E\mathfrak{f})}\,.$ (79)

### 23 Jan 2024 v3

@@ 1 Introduction
While quantum heat engines have been extensively studied within the domain of quantum mechanics, their exploration in the context of relativistic quantum mechanics Munoz.QHE.relativistic.dirac.2012 ; Pena.RQM.engine.2016 ; Purwanto.RQHE.2016 ; YIN201858 ; Chattopadhyay.relativistic.uncertainty.2019 ; Myers.QOE.relativistic.2021 ; Chattopadhyay.noncommutative.thermo.2021 ; Sukamto.RQHE.minimal.length.2023 and quantum field theory remains relatively unexplored. Recent investigations UnruhOttoEngine ; Finn.UnruhOtto have delved into QOEs where the working substance is an Unruh-DeWitt (UDW) particle detector Unruh1979evaporation ; DeWitt1979 interacting with quantum scalar and spinor fields. In their QOE, the field is assumed to be in the Minkowski vacuum state, and work is extracted from a thermal bath induced by the Unruh effect Unruh1979evaporation experienced by a linearly accelerating detector. Such a QOE driven by the Unruh effect is particularly called the Unruh QOE. See also Xu.UnruhOtto.degenerate ; Kane.entangled.Unruh.Otto ; Barman.entangled.UnruhOtto ; Mukherjee.UnruhOtto.boundary for subsequent studies. The Unruh QOE concept extends to scenarios where a UDW detector follows an arbitrary timelike trajectory in curved spacetime Gallock2023Otto . We will refer to quantum heat engines that include relativistic effects as ‘relativistic quantum heat engines.’
@@ 3.1 Algebraic quantum field theory
The CCR algebra $\mathcal{A}(\mathcal{M})$ is generated by smeared field $\phi(f)$ obeying the following relations khavkine2015algebraic , which we refer to as the Klein-Gordon relations:
@@
Let us define a ***-representation of $\mathcal{A}(\mathcal{M})$ as a triple $(\mathcal{H},\mathcal{D},\pi)$, where $\mathcal{H}$ is Hilbert space, $\mathcal{D}\subset\mathcal{H}$ is a dense subspace of $\mathcal{H}$, and $\pi$ (also referred to as a representation) is a linear homomorphism,
@@ 5.2 Observations and remarks
    The positive work condition (55) implies that if the detector is unable to self-communicate, then $E(\mathfrak{f}_{1},\mathfrak{f}_{2})=0$ leading to $\braket{W_{\text{ext}}}=0$. Thus, self-communication is necessary to extract positive work for a delta-coupled UDW detector. We stress that this holds for any globally hyperbolic spacetimes, though the way signals propagate may vary depending on the spacetime under consideration. It should be noted that merely receiving a signal from the past does not guarantee work extraction since the sign in (55) might be positive even when $E(\mathfrak{f}_{1},\mathfrak{f}_{2})\neq 0$. We will address the mechanism for the work extraction at the end of this subsection.
@@
The ranges of $p_{2}$ in (58) are due to the fact that $\nu_{\mathfrak{f}_{1}}\nu_{\mathfrak{f}_{2}}\alpha\in(0,1]$, as given by (51).
@@ 5.3 Example: An inertial detector extracts work from the Minkowski vacuum
Figure 5: The extracted work $\braket{W_{\text{ext}}}\sigma$ in units of the characteristic width $\sigma$ of the Gaussian smearing function. (a) Work as a function of the second interaction time $\tau_{2}/\sigma$ when $\tau_{1}/\sigma=0$, $\Omega_{1}\sigma=1$, $\Omega_{2}\sigma=3$, $\tilde{\lambda}_{1}/\sigma=100$, and $\tilde{\lambda}_{2}/\sigma=1$. Negative $\braket{W_{\text{ext}}}\sigma$ corresponds to work done on the detetor. The yellow plot shows the positive work condition (55). (b) Work as a function of the coupling constants. Here, we choose $\tau_{1}/\sigma=0$, $\Omega_{1}\sigma=1$, $\Omega_{2}\sigma=3$, and $\tau_{2}/\sigma=1.5$.
@@
After substituting these expressions into the extracted work (54), the results are shown in Figure 5. In Figure 5(a), we depict the dependence of the second interaction’s timing, $\tau_{2}/\sigma$, in units of the characteristic Gaussian width $\sigma$, given $\tau_{1}/\sigma=0$, $\Omega_{1}\sigma=1$, $\Omega_{2}\sigma=3$, $\tilde{\lambda}_{1}/\sigma=100$, and $\tilde{\lambda}_{2}/\sigma=1$. As discussed in Section 5.2, the detector can extract positive work from the vacuum state of the field instantaneously as long as the first interaction affects the second one. In fact, as one can see from the positive work condition depicted as the yellow curve in 5(a), we have $\braket{W_{\text{ext}}}\sigma=0$ once the detector cannot receive the signal from the first interaction (i.e., when $\tau_{2}/\sigma>4$).
@@ 6 Conclusion
Note: A related study was recently conducted in kollas2023exactly . In this study, the authors consider a relativistic QOE utilizing a delta-coupled UDW detector in Minkowski spacetime to extract work from a quantum field in a thermal state.
@@ Appendix A Second interaction
$\nu_{\mathfrak{f}}$ $=\omega(W_{2\mathfrak{f}})=e^{-2\mu(E\mathfrak{f},E\mathfrak{f})}\,.$ (79)